Calculate:
1. CPU per basket
2. The inflation rate
If in the CPU basket contain the quantities bought in 2010. The average household spending $60 on juice and $30 on cloth in 2009.When the price of the juice was $2 a bottle and the price of cloth was $5 meter.In the current year 2010 juice is $4 a bottle and cloth is $6 a meter.
1
Expert's answer
2019-08-14T09:25:09-0400

1. CPU in 2009 is 100.

CPI in 2010 is:

CPI = (60/2×4 + 30/5×6)/(60 + 30)×100 = 173.33.

2. The inflation rate in 2010 is: (173.33 - 100)/100×100% = 73.33%.